OpenStudy (anonymous):

which polynomial is not a factor of 3x^4-10x^3-29x^2+16x+20?

OpenStudy (anonymous):

A)x-1 B)3x+2 C)x+2 D) x-2 I think it is A)

OpenStudy (whpalmer4):

If the polynomial is a factor of the big polynomial, then setting the little polynomial equal to 0 and solving for \(x\) will yield a root or zero of the big polynomial. A zero (or root) has the characteristic that the polynomial's value when evaluated at x = root is identically 0. So, if x-1 is a factor, x-1=x, x = 1, and \[3(1)^4-10(1)^3-29(1)^2+16(1)+20 = 0\] If that is true, then \(x-1\) is a factor of the polynomial \(3x^4-10x^3-29x^2+16x+20\)

OpenStudy (whpalmer4):

sorry, that should have been "if x-1 is a factor, x-1=0, x = 1"